The Bob of a pendulum has a maximum velocity of 2m/s at it's lowest position, O. calculate the height h of the Bob above O, where it's velocity is zero (Take g=10m/s²)


1
Expert's answer
2021-04-05T11:08:53-0400

We can find the height "h" of the Bob above O from the law of conservation of energy:


"KE=PE,""\dfrac{1}{2}mv^2=mgh,""h=\dfrac{v^2}{2g}=\dfrac{(2\ \dfrac{m}{s})^2}{2\cdot10\ \dfrac{m}{s^2}}=0.2\ m."